OpenStudy (anonymous):

find the inverse of f(x)=5x-4

OpenStudy (anonymous):

y=5x-4 x=5y-4 x-4=5y (x-4)/5=y f(x)=(x-4)/5

OpenStudy (anonymous):

I'm now sure with mine too :D I was solving that questions when i was 14-15 :D

OpenStudy (anonymous):

But i guess lamel1993 didn't mean multiplicative inverse